OpenStudy (anonymous):

find all imaginary solutions of x^4+14x^2-32 by factoring

sam (.sam.):

Cubed will give 2 reals 2 im's

OpenStudy (anonymous):

do u know how'd start?

sam (.sam.):

I mean to the forth power

OpenStudy (anonymous):

factor the common variable out

sam (.sam.):

Great, what do you get?

sam (.sam.):

We can set \[f(x)=x^4+14x^2-32\] When we're finding the zeros, f(x) will be zero because f(x) is y, and when x-intercepts, y will be zero, so \[f(x)=0\] ----------------------------------- \[x^4+14x^2-32=0\] \[\left(x^2-2\right) \left(x^2+16\right)=0\]

OpenStudy (anonymous):

thank you but it says your suppose to have 4 answers that equal x because the leading term is to the fourth degree

sam (.sam.):

Yes, you will have 4 answers, \[x^2-2=0~~~~~and~~~~~x^2+16=0 \\ \\ x=\pm \sqrt{2} ~~~~~and~~~~~ x= \pm \sqrt{-16}\] ---------------------------------------------- \[x=\sqrt2 ~~~~~ x=-\sqrt{2}~~~~~~~~~|~~~~~~~~x=4i~~~~~x=-4i\]

OpenStudy (anonymous):

I think that is how you do it. thank you. but did you plug in x^3 an x in the equation to get those answers?